Steven walks 11 feet to the East, 18 feet to the west, and 20 feetA. What is the total distance traveled?B. What is Steve’s total displacement?Pleasee helpp!!

1
Expert's answer
2021-01-20T08:43:32-0500

a)


"d=11+18+20=49\ ft"

b)


"s=11-18+20=13\ ft"

Direction: to the east.